# The Fine-Tuning Paradox: Boosting Translation Quality Without Sacrificing LLM Abilities

### v1

@@ 1 Introduction
  • We introduce a novel evaluation dataset, IdiomsInCtx-MT, to measure non-literalness performance. To our knowledge, it is the first dataset that consists of idiomatic expressions in context and their human-written translations. It covers 2 language pairs with 3 translation directions.
@@ 2 Related Work
There are multiple strategies for fine-tuning LLMs for machine translation. One approach makes use of either a small set of high-quality human-written translations or a set of translation instructions for fine-tuning (Li et al., 2023; Zeng et al., 2023; Jiao et al., 2023). Another line of work makes use of more traditional machine translation data: parallel data from the web, which is orders of magnitude larger compared to what is used in fine-tuning (Yang et al., 2023; Alves et al., 2023; Zhang et al., 2023; Zhu et al., 2023b). These strategies are focused on improving general machine translation quality, but it remains unclear what happens to other abilities that are relevant for translation. We investigate the effect of fine-tuning on relevant abilities for translation using publicly available models.
@@ 3.1 Experimental Setup
Following Xu et al. (2023), we use human-written translations from WMT17 to WMT20, resulting in 89K training examples that are evenly distributed across the language directions we consider. On this dataset, we perform full fine-tuning for models up to 40B and QLoRA fine-tuning for the LLaMA 65B model.
@@
We use a subset of 1.4M sentence pairs of this filtered data that is evenly distributed across language directions for training. We fine-tune LLaMA models up to 40B parameters on this dataset but leave out larger models because of the high computational cost.
@@
  • WMT22 To evaluate general machine translation quality, we use WMT22 (Kocmi et al., 2022) test sets consisting of news, e-commerce, social, and conversational domains. We evaluate all language directions on this test set in a 0-shot setting. Following the recommendation of Kocmi et al. (2021), we report COMET scores111model: unbabel/wmt22-comet-da (Rei et al., 2020). We do not report BLEU scores (Papineni et al., 2002), since this metric is known to have a poor correlation with human judgments (Mathur et al., 2020; Freitag et al., 2021; Kocmi et al., 2021; Freitag et al., 2022).
  • CoCoA-MT To evaluate formality steering ability of LLMs, we make use of the CoCoA-MT (Nadejde et al., 2022) dataset. It consists of 600 test sentences with English source and contrastive target sentences consisting of a formal and informal translation. We use German as the target language and report both COMET and accuracy based on the ratio of correctly predicted formality forms. We use 5-shot examples to bias the formality of outputs. [^2] We also experimented with steering formality through prompting, but the results were inferior to using 5-shot examples.
@@
  • ctxpro We evaluate 0-shot performance on longer inputs that includes sentences that require context to be disambiguated correctly by including ctxpro (Wicks and Post, 2023). We consider the animacy ambiguity type in German $\rightarrow$ English and Russian $\rightarrow$ English.  [^3] We also experimented with different ambiguity types (auxiliary and gender in out-of-English directions). However, the resulting translations were often incomplete, containing only a subset of the total number of sentences, making it impossible to effectively evaluate decontextualization abilities. While English makes no gender distinction for inanimate objects, some other languages such as Russian do. The ambiguous animacy examples in the ctxpro dataset require corresponding document-level context for correct disambiguation. We subsample the test set to 2K examples per language direction. The average number of sentences per input is $10.16\pm 1.27$. We report the generative accuracy score, which measures the accuracy of the contextualization.
@@
Additionally, we considered style transfer and constrained generation, such as prompting a language model to use custom terminology, as additional translation capabilities. However, our final choices were influenced by the availability of suitable test sets and the applicability of targeted automatic metrics, leading us to not pursue these options.
@@ 3.2 Results
We show results for formality steering using 5-shot examples for models trained on human-written data in Figure 2. Notably, the base models exhibit strong performance for this task; for instance, the LLaMA-7b model achieves an accuracy of 0.982 in identifying informal markers. However, fine-tuning on only 18K examples results in a decline of this ability to 0.862, even though German-English COMET on WMT22 continues to improve up to 36K examples. The degradation is more pronounced with informal markers, which is likely attributable to the formal bias inherent in the WMT22 training data. Fine-tuning on more data further degrades formality steering capabilities: there is a significant negative correlation (Spearman’s $\rho=-0.46$, $p<0.001$) between formal and informal marker prediction and dataset size. We find that COMET stays relatively constant (see Appendix B Figure 8), despite some fluctuations, indicating that it does not adequately capture formality markers. As a result, the accuracy scores correlate very weakly ($\rho=0.16$, $p<0.001$) with COMET scores, which suggests that comprehensive evaluation of LLMs for machine translation benefits from task-specific metrics.
Further, when examining models trained on OPUS data (see Appendix B Figure 9), we find that increasing the amount of fine-tuning data up to 1.4M parallel sentences further exacerbates the degradation. Again, we observe a significant negative correlation ($\rho=-0.58$, $p<0.001$) between accuracy and dataset size for OPUS-trained models, reinforcing the conclusion that larger parallel datasets during fine-tuning adversely affect formality steering capabilities.
@@
Similar to the WMT22 test set, we see that COMET scores improve until the first 1-3 checkpoints before stabilizing or starting to decrease. However, for idiomatic expressions even thed targeted metrics LitTER and MWEScore improve during fine-tuning or least remain stable. This indicates that even a large open-source model like LLaMA-65b still has room for improvement when it comes to idiomatic translations. Future work could investigate translation literalness and idiomaticity of LLM translations on stronger models such as GPT-3.5 during fine-tuning. [^5] https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates
@@ 5 Conclusion
Because of the high cost of repeatedly fine-tuning LLMs of different sizes, we limited ourselves to 6 translation directions (German, Chinese, and Russian in and out of English). The impact of fine-tuning on emergent abilities of LLMs when translating in and out of low-resource languages is not studied in our work. It is therefore possible that our findings do not generalize to low-resource languages.
@@ B.2 Formality steering
Figure 8 shows COMET scores for models trained on WMT data. COMET scores stay relatively constant, in contrast to the accuracy of formality forms.
@@
Figure 8: COMET scores on CoCoA-MT for models trained on WMT data.

### v2

@@ 1 Introduction
  • We introduce a novel evaluation dataset, IdiomsInCtx-MT [^1] We release the dataset at https://github.com/amazon-science/idioms-incontext-mt, to measure non-literalness performance. To our knowledge, it is the first dataset that consists of idiomatic expressions in context and their human-written translations. It covers 2 language pairs with 3 translation directions.
@@ 2 Related Work
There are multiple strategies for fine-tuning LLMs for machine translation. One approach makes use of either a small set of high-quality human-written translations or a set of translation instructions for fine-tuning (Li et al., 2023; Zeng et al., 2023; Jiao et al., 2023; Wu et al., 2024; Zhu et al., 2024). Another line of work makes use of more traditional machine translation data: parallel data from the web, which is orders of magnitude larger compared to what is used in fine-tuning (Yang et al., 2023; Alves et al., 2023; Zhang et al., 2023; Zhu et al., 2023b). These strategies are focused on improving general machine translation quality, but it remains unclear what happens to other abilities that are relevant for translation. We investigate the effect of fine-tuning on relevant abilities for translation using publicly available models.
@@ 3.1 Experimental Setup
Following Xu et al. (2023), we use human-written translations from WMT17 to WMT20, resulting in 89K training examples that are evenly distributed across the language directions we consider.
@@
We use a subset of 1.4M sentence pairs of this filtered data that is evenly distributed across language directions for training. We fine-tune LLaMA models up to 30B parameters on this dataset but leave out larger models because of the high computational cost.
@@
  • WMT22 To evaluate general machine translation quality, we use WMT22 (Kocmi et al., 2022) test sets consisting of news, e-commerce, social, and conversational domains. We evaluate all language directions on this test set. We evaluate in a 0-shot setting, and report COMET222model: unbabel/wmt22-comet-da (Rei et al., 2020) scores.
  • CoCoA-MT To evaluate formality steering ability of LLMs, we make use of the CoCoA-MT (Nadejde et al., 2022) dataset. It consists of 600 test sentences with English source and contrastive target sentences consisting of a formal and informal translation. We use German as the target language and report accuracy based on the ratio of correctly predicted formality forms. We use 5-shot examples to bias the formality of outputs. [^3] We also experimented with steering formality through prompting, but the results were inferior to using 5-shot examples.
@@
  • ctxpro We evaluate performance on longer inputs that includes sentences that require context to be disambiguated correctly by including ctxpro (Wicks and Post, 2023). We consider the animacy ambiguity type in German $\rightarrow$ English and Russian $\rightarrow$ English. [^4] We also experimented with different ambiguity types (auxiliary and gender in out-of-English directions) but the resulting translations are of poor quality, making it impossible to sensibly evaluate decontextualization abilities. While English makes no gender distinction for inanimate objects, some other languages such as Russian do. The ambiguous animacy examples in the ctxpro dataset require corresponding document-level context for correct disambiguation. We subsample the test set to 2K examples per language direction. The average number of sentences per input is $10.16\pm 1.27$. We report the generative accuracy score, which measures the accuracy of the contextualization.
@@ 3.2 Results
We show results for formality steering using 5-shot examples for models trained on human-written data in Figure 2. Notably, the base models exhibit strong performance for this task; for instance, the LLaMA-7b model achieves an accuracy of 0.982 in identifying informal markers. However, fine-tuning on only 18K examples results in a decline of this ability to 0.862, even though German-English COMET on WMT22 continues to improve up to 36K examples. The degradation is more pronounced with informal markers, which is likely attributable to the formal bias inherent in the WMT22 training data. Fine-tuning on more data further degrades formality steering capabilities: there is a significant negative correlation (Spearman’s $\rho=-0.46$, $p<0.001$) between formal and informal marker prediction and dataset size. Notably, these accuracy scores correlate very weakly ($\rho=0.16$, $p<0.001$) with COMET scores, which suggests that comprehensive evaluation of LLMs for machine translation benefits from task-specific metrics. Further, when examining models trained on OPUS data (see Appendix B Figure 8), we find that increasing the amount of fine-tuning data up to 1.4M parallel sentences further exacerbates the degradation. Again, we observe a significant negative correlation ($\rho=-0.58$, $p<0.001$) between accuracy and dataset size for OPUS-trained models, reinforcing the conclusion that larger parallel datasets during fine-tuning adversely affect formality steering capabilities.
@@
Similar to the WMT22 test set, we see that COMET scores improve until the first 1-3 checkpoints before stabilizing or starting to decrease. However, for idiomatic expressions even the targeted metrics LitTER and MWEScore improve during fine-tuning or least remain stable. This indicates that even a large open-source model like LLaMA-65b still has room for improvement when it comes to idiomatic translations. Future work could investigate translation literalness and idiomaticity of LLM translations on stronger models such as GPT-3.5 during fine-tuning. [^5] https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates
@@ 5 Conclusion
Because of the high cost of repeatedly fine-tuning LLMs of different sizes, we limited ourselves to 6 translation directions (German, Chinese, and Russian in and out of English). The impact of fine-tuning on emergent abilities of LLMs when translating in and out of low-resource languages is not studied in our work.